If the sum of a fathers and sons age is 53, and the son is 25 years younger then the dad how old is the son
Answer by math-vortex(648)

If the sum of a fathers and sons age is 53, and the son is 25 years younger then the dad how old is the son.
==================
Let
f = age of the father
.
If the son is 25 years younger than the father, then an algebraic expression for the son's age is f - 25.
.
We know that the father's age and the son's age add up to 53. In algebra, we write that as
f + (f - 25) = 53
.
Now we just solve this equation for f.
2f - 25 = 53
2f = 53 + 25
2f = 78
f = 78/2
f = 39
.
So the father is 39 years old. Since the son is 25 years younger, the son is 39 - 25 = 14 years old.
.
Checking, we see that their ages add up to 53 because 14 + 39 = 53.
Bingo!
.
Hope this helps!
.
Ms.Figgy
math.in.the.vortex